8 years ago a dad was 8 times older than his son.
Today their combined ages are 52, how old are they today?

Answer by Maths68(1474)

Let
Present age of Dad = d
Present age of Son = s

Today their combined ages are 52
d+s=52
d=52-s.........(1)
8 years ago a dad was 8 times older than his son.
d-8=8(s-8)
d-8=8s-64
d=8s-64+8
d=8s-56.......(2)
Put the value of d from (1) to (2)
52-s=8s-56
-s-8s=-56-52
-9s=-108
-9s/-9=-108/-9
s=12
Put the value of s in (1)
d=52-12
d=40
Present age of Dad = d = 40 years old
Present age of Son = s = 12 years old